find the measure of a supplement of angle R where angle R is equal to 6x+15
Answer by nerdybill(6959)

find the measure of a supplement of angle R where angle R is equal to 6x+15
.
If two angles are supplementary, the sum of the two angles = 180
.
Therefore, the
"supplement of angle R" = 180 - (6x+15)
"supplement of angle R" = 180 - 6x -15
"supplement of angle R" = 165 - 6x